OpenStudy (anonymous):

Kittens weigh about 100 grams when born and gain 7 to 15 grams per day. If a kitten weighed 100 grams at birth and gained 8 pounds per day, in how many days will the kitten triple it's weight? How do I write this as an equation, and how do I find the answer?

OpenStudy (anonymous):

Hmmm... make it into an equation.

OpenStudy (anonymous):

how would i do that?

OpenStudy (anonymous):

8x25=200 200+100=300 100x3=300 It will take 25 days for the cat to reach triple its original weight.

OpenStudy (anonymous):

can you explain this, please?

OpenStudy (anonymous):

i need it in an algebraic equation, with a variable

OpenStudy (anonymous):

300=100+8x

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

Now to further explain. 300=100+8x (I have to get x alone) 300-100=100-100+8x (i move the 100 over to the other side) 200=8x (now i divide both sides by 8) 200/8=8/8x (the 2 8's cancel and 8 divides evenly into 200) 25=x

OpenStudy (anonymous):

you... are... SO SMART!!!!! and thanks to you, now i'll be smart, too. thanks again. :) :)

OpenStudy (anonymous):

Good luck and you're welcome

OpenStudy (anonymous):

Sorry, Pika. So many people >.<

OpenStudy (anonymous):

its ok :) PreCalTroubles solved my troubles :) thanks anyway :)